OpenStudy (anonymous):

exact value of: csc (11pi)/4

OpenStudy (btaylor):

Are you looking for \[\frac{\csc(11\pi)}{4}\] or \[\csc(\frac{11 \pi}{4})\]?

OpenStudy (anonymous):

find \(\frac{11\pi}{4}\) on the unit circle on the last page of the attached cheat sheet then look at the coordinates the second coordinate is sine, the cosecant is the reciprocal of that

OpenStudy (anonymous):

actually that is not much help i see because you will not find \(\frac{11\pi}{4}\) but if you subtract \(\frac{11\pi}{4}-2\pi \) you get \(\frac{3\pi}{4}\) which is coterminal, use \(\frac{3\pi}{4}\) instead